### v1

# Engineering Carbon Credits Towards A Responsible FinTech Era: The Practices, Implications, and Future

Abstract
Carbon emissions are a key factor to serious climate change, and carbon credits have emerged as an effective tool for mitigating climate deterioration and environmental damage while assisting organisations in managing their carbon footprint. In recent years, carbon credits have garnered significant attention across various economic sectors due to their substantial impact. However, fully utilizing them to address these issues remains a challenge. This study aims to enhance understanding of the engineering practices for carbon credits to develop responsible fintech solutions, thereby providing stronger incentives and insights for carbon emission management. To achieve this goal, we first conduct a comprehensive review of the negative impacts associated with organisations’ strategy of evading carbon management through non-disclosure of carbon emissions. Our findings reveal that both non-disclosure of carbon emissions and high carbon emissions negatively affect a organisation’s financial stability and market value. Therefore, it is suggested that organisations should actively manage their carbon emissions and transparently share relevant information to mitigate potential risks. This study highlights the importance of carbon credits for businesses, emphasizing their role in improving corporate sustainability. With the insights from the carbon emission, we then investigate the engineering methods that assist organisations in managing carbon emissions more cost-effectively, in particular for the data-driven computing solutions. Specifically, we analyse the factors influencing carbon prices and review the state-of-the-art carbon price prediction algorithms. These tools help organisations optimize their carbon credit purchasing strategies, thereby reducing costs and improving operational efficiency. Furthermore, we review the most recent corporate carbon emission prediction algorithms, which not only provide more accurate environmental performance assessments for investors and governments when carbon emissions data is not disclosed, but also help organisations estimate their future carbon credits needs, supporting better budget planning and strategy optimization. Finally, by integrating carbon price and carbon emission predictions, this paper proposes future research directions, including the prediction of corporate-level carbon management costs. We anticipate this will lay a foundation for future quantitative research on how carbon management practices impact corporate market value and financial performance. To the best of our knowledge, this is the first systematic review covering the various aspects of carbon credits, with a particular focus on computing solutions and engineering practices.
1 Introduction
In today’s global landscape, the Industrial Revolution mark a turning point where fossil fuels become the primary energy source, driving economic growth but also significantly increasing carbon emissions [1]. This rise in carbon emissions has led to serious climate issues, as excessive carbon dioxide forms a layer around the Earth’s surface. This layer traps more heat while reducing its dissipation, directly leading to the global temperature increases. The resulting climate change is profound and often irreversible effects on ecosystems, including glacier melting and biodiversity loss. Additionally, rising sea levels also pose a grave threat to human settlements, particularly in coastal regions. In response, the global community has invested substantial resources in promoting climate-friendly solutions [2]. A key initiative, the Paris Agreement, reflects the collective commitment of nations to limit global temperature increases to within 2 degrees Celsius by the end of this century [3].
The Carbon Emissions Trading (CET) system, proposed under the Kyoto Protocol, is a key strategy to combat climate change. Introduced in 1997, CET aims to reduce greenhouse gas emissions by implementing a ‘quota trading’ mechanism. Each unit of carbon credit allows the emission of one ton of carbon dioxide [4]. Specific emission quotas are allocated to countries based on their respective development status and goals, which is then distributed to domestic organizations and organisations. Entities that exceed their allocated emissions must purchase carbon credits on the carbon trading market, while those with surplus credits can sell them on the market. This scheme has been adopted in many regions, including Europe, Asia, and the Americas [5]. In addition to regulating global carbon emissions, CET incentivises innovation in green technologies and promotes the use of renewable energy sources to further reduce carbon emissions.
As the carbon credit mechanism matures, it role in mitigating global temperature rise is increasingly recognised. Decision-making must now integrate strategies for managing both carbon emissions and the carbon credits cost. For organisations with high emissions, the cost of purchasing additional carbon credits can be substantial [6], making accurate carbon credit price prediction essential for effective cost management [7].
Additionally, many organisations opt to either inaccurately disclose their carbon emissions or avoid voluntarily disclosing their actual carbon emissions in practise [8]. This creates significant risks for businesses, particularly when managing carbon emission responsibilities that are closely tied to corporate reputation and financial performance. For example, some organisations fail to actively manage their carbon emissions, particularly by not purchasing carbon credits to offset excess emissions. Such practices not only clearly violate policy regulations but also expose organisations to greater compliance risks. To avoid the potential damage to their corporate image, such as fines, a decrease in market value, and harm to their reputation, many organisations choose not to disclose their actual carbon emissions, carbon management practices, or the associated operational costs in their financial reports [9]. In this paper, our findings reveal that better engineering the carbon credits for industry sectors can enable an enhanced transparency of the process, which can significantly boost the organisation performance.
There are two key reasons for the increased transparency in carbon emissions disclosure. First, in the financial markets, increasing attention is being paid to whether organisations adhere to ESG (Environmental, Social, and Governance) principles. Investors, particularly institutional investors, are placing greater emphasis on organisations’ performance in sustainable development [10]. They are more inclined to allocate funds to organisations that can create long-term social value and mitigate negative externalities. This trend has significantly heightened the demand for ESG information disclosure, with investors favoring organisations that actively implement ESG principles. Consequently, financial investors typically rely on the information disclosed in organisations’ financial reports to assess their ESG performance. However, some organisations have not fully disclosed their carbon emissions and carbon management practices, making it difficult for investors to gain a comprehensive understanding of the company’s actual operations and true value. This information asymmetry not only increases uncertainty and risk in investment decisions but also potentially leads investors to miss identifying high-risk organisations. Therefore, having accurate and comprehensive information on a company’s carbon management is crucial for investors to make informed decisions.
Second, government has implemented stricter regulations on organisations’ disclosure of carbon emissions. Governments also need to understand a company’s carbon emissions and management practices to formulate more effective carbon management policies and enforce penalties on non-compliant organisations, thereby supporting the achievement of emission reduction targets and controlling global temperature rise. By obtaining accurate carbon emission data, governments can better evaluate the effectiveness of current policies and adjust them in a timely manner to ensure that carbon emissions align with expected targets [11]. Additionally, understanding a company’s carbon management strategies helps governments identify industries or organisations that are lagging in emission reduction efforts, enabling the formulation of targeted support or punitive measures to promote the overall industry’s green transformation. Furthermore, governments can use this data in international climate negotiations to showcase the country’s efforts and achievements in emission reductions, thereby strengthening international confidence in the country’s climate commitments. To address the challenges faced by organisations, financial investors, and governments as outlined above, we investigate following three areas:
  1. 1.
    The impact of non-disclosure of carbon emissions and carbon management information on organisations. This study explores whether non-disclosure can effectively help organisations avoid penalties and negative consequences when they fail to manage carbon emissions as required.
    • Findings: This study finds that organisations that fail to disclose information about their carbon emissions and management not only risk negative impacts but may also face even more severe consequences. Therefore, evasion and non-disclosure are unsustainable strategies. Instead, organisations need to proactively manage carbon emissions, fulfil their social responsibilities, and transparently disclose relevant information. This approach not only helps organisations enhance regulatory compliance and market trust, but also allows them to gain a competitive advantage in markets that increasingly prioritize ESG principles.
  2. 2.
    An analysis of the latest carbon credit price prediction algorithms and factors influencing carbon prices.
    • Findings: This study systematically reviews and analyses the state-of-the-art methods for predicting carbon credit prices. These methods help organisations accurately identify optimal times to purchase carbon credits when prices are low, thereby reducing carbon management costs and providing economic incentives for optimizing carbon management. By lowering the cost of purchasing carbon credits, organisations can reduce expenditures while boosting profits, which in turn encourages them to voluntarily and transparently disclose their carbon emissions. This process not only helps organisations improve their compliance and social responsibility but also strengthens the trust of investors and regulators in the company’s carbon management efforts, fostering a virtuous cycle of proactive corporate participation in the carbon market.
  3. 3.
    An analysis of the latest corporate carbon emission prediction algorithms.
    • Findings: This study systematically reviews the most effective corporate carbon emission prediction methods available. These methods provide accurate emission assessments even when organisations fail to disclose or misreport carbon emission data. Through these prediction methods, financial investors can more precisely evaluate an organisation’s environmental performance and long-term value, leading to more informed investment decisions. Simultaneously, governments can use these data to develop more effective carbon emission policies, ensuring that regulatory measures are implemented and enforced. Additionally, these methods enable governments to identify and monitor an organisation’s actual carbon emissions, allowing for appropriate management and penalties when necessary. This approach not only reinforces corporate accountability but also enhances market transparency, ultimately contributing to the achievement of global emission reduction targets.
Although the existing literature increasingly focuses on the application of carbon credits, there lacks comprehensive reviews on this topic. Existing reviews tend to be limited on a specific area [2, 12, 13, 14], such as the impact of carbon credits on farmers’ income in agricultural, or the role of blockchain for carbon credits in the construction industry. However, these studies do not provide a systematic review of carbon credits from a financial and fintech perspective, leaving the broader implication for enterprises, governments, and the financial industry largely unexplored. In particular, there is a significant gap in understanding and leveraging carbon credits in fintech with advanced computing solutions.
This study addresses the existing gap by exploring carbon emission and carbon price prediction technologies, providing stakeholders with more comprehensive management tools. Following the findings of carbon credit price and carbon emission prediction algorithms, we present a thorough understanding of the significance of carbon credits, in particular for the organisation performance. We also identify a promising yet challenging task as the future research direction to combine these two important tasks to further enhance the carbon management practices, which will in turn increase the corresponding market value.
Thus, in this work, we conduct a systematic literature review [15], to comprehensively analyse the application of carbon credits in fintech sector, providing valuable insights for stakeholders in tackling climate change challenges with state-of-the-art computing algorithms and solutions. The main contributions of this article are as follows:
  1. 1.
    This study systematically summarizes the impact of transparency in carbon emission management on corporate development. It explores the strategies that organisations use to mitigate compliance risks by not disclosing carbon emission information, analysing the effectiveness and potential consequences of such approaches. The research shows that a lack of transparency not only fails to protect organisations from negative impacts of policies and markets but may also exacerbate the risks they face. Following sections highlight the current best practices of distilling carbon emission management in a measurable approach, in which we identify the topics of estimations of carbon credit price and corporate carbon emission.
  2. 2.
    This study provides a comprehensive review of the factors influencing carbon credit prices, offering insights that can qualitatively predict price trends. This contribution aids organisations in better managing their carbon credit purchasing strategies, thereby reducing costs. Additionally, these factors can serve as input variables for algorithms predicting carbon price fluctuations, further enhancing the accuracy and applicability of these algorithms in supporting corporate decision-making in the carbon market.
  3. 3.
    The article systematically summarizes and critically evaluates the applicability, limitations, and advantages of current carbon credit price prediction algorithms and methods. This analysis supports organisations in making informed decisions about the timing and scale of carbon credit purchases, thereby promoting more effective carbon management practices.
  4. 4.
    The article reviews and discusses the latest corporate carbon emission prediction algorithms, highlighting their strengths and weaknesses. This contribution assists investors and regulators in accurately assessing a company’s environmental performance, even in cases of non-disclosure, thereby reducing information asymmetry and enhancing market transparency.
  5. 5.
    This study, through an in-depth exploration of carbon emission and carbon price prediction technologies, reveals the significant impact of carbon credits on corporate market value. These technologies not only help organisations strategically plan their carbon credit purchases and reduce carbon management costs, but more importantly, they provide a foundation for quantitative research, enabling organisations to predict the impact of their carbon management practices on financial and market performance. By actively managing carbon emissions and transparently disclosing related information, organisations can enhance market trust, reduce financing costs, and ultimately increase their market value. This lays a strong foundation for future research to quantify the long-term impact of carbon management on corporate market value.
The remaining structure of the article is as follows: Section 2 provides a detailed description of the procedure used to select papers for this systematic review, including the inclusion and exclusion criteria for literature retrieval. Section 3 discusses the impact of corporate transparency in carbon emission disclosure. Section 4 conducts an in-depth analysis of the key factors influencing the price of carbon credits. Section 5 examines algorithms related to predicting carbon credit prices, focusing on their applicability and effectiveness. Section 6 explores methods for predicting corporate carbon emissions, discussing the strengths and weaknesses of various algorithms. Section 7 provides prospects for the development of carbon credits in the financial and fintech fields. Section 8 compares this study with relevant surveys and reviews. Section 9 summarizes the key findings and contributions of this study.
2 Research methodology
This study adopts a systematic literature review approach [16] to comprehensively identify and analyse the key applications of carbon credits in the fin-tech field, as well as the research progress in related areas. To ensure the transparency and replicability of this work, we strictly follow the steps, including initial literature retrieval, setting up inclusion and exclusion criteria, performing study selection process and finally conducting quality assessment.
To begin with, we consider below the key research questions for investigation:
  1. 1.
    What is the impact of corporate carbon emissions disclosure (or non-disclosure) on organisations and society, and what factors influence corporate decisions regarding such disclosures? (Section. 3)
  2. 2.
    What are the key factors that influence carbon credit prices? (Section. 4)
  3. 3.
    What methods are available for predicting carbon prices? (Section. 5)
  4. 4.
    What methods are available for forecasting corporate carbon emissions? (Section. 6)
2.1 Step 1: File Retrieval
In this first step, we identified the key data sources and search terms used in the study. For the data sources, they are generally from Google scholar, Scopus, IEEE Xplore, and Science Direct. The results are collected in August 2023 and include articles from 2005 to 2024. This timeframe is chosen because key terms such as ‘carbon credits’ and ‘carbon emissions’ emerged around 2005. The search strings were constructed using key phrases from our research questions, and the search terms were combined using ‘AND’ and ‘OR’ logical operators to ensure a comprehensive search.
The selected databases are among the most widely used academic sources globally, renowned for their comprehensive and authoritative coverage. We reviewed past research papers to understand the current status and gaps in carbon credit systems, focusing on factors influencing carbon prices, algorithms for predicting carbon credit prices, algorithms for forecasting enterprise emissions, and the impact of carbon emissions disclosure on organisations.
2.2 Step 2: Inclusion and Exclusion Criteria
Herein we provide our inclusion and exclusion criteria for literature. For the inclusion criteria:
  1. 1.
    Study Types: Original research articles, conference papers, and case studies that provide new insights or data analysis to enhance understanding of the carbon credit market, price dynamics, or influencing factors.
  2. 2.
    Publication Time Frame: Considering the development of the carbon market and related policies, only literature from 2005 to present is considered to ensure relevance and timeliness of the research.
  3. 3.
    Research Focus: Literature must focus on carbon credits, carbon pricing strategies, market dynamics analysis, policy impact assessments, or the application of carbon credits in fintech.
  4. 4.
    Language: Only English-language literature is considered.
Exclusion Criteria:
  1. 1.
    Non-Peer-Reviewed Literature: Exclude literature that has not undergone peer review, such as news articles, blog posts, and informal white papers, to ensure the scientific validity and authority of the research.
  2. 2.
    Case Study Limitations: While case studies are included, those based solely on a single case and fail to provide generalised insights or in-depth analysis will be excluded.
  3. 3.
    Geographic Scope: Exclude studies focused on specific regions with known extreme immaturity of the carbon credit market or inconsistent with global trends in carbon market development.
  4. 4.
    Outdated Literature: While studies since 2005 are considered, those based on outdated or no longer applicable data, models, or assumptions are excluded.
  5. 5.
    Professional Books and Chapters: Exclude entire books and book chapters, as these are usually not peer-reviewed and are less accessible and timely than journal articles and conference papers.
  6. 6.
    Review Articles and Commentaries: Exclude review articles and commentaries that do not provide original research or new data, unless they offer unique insights into specific issues or integrate the latest developments from a large body of research.
  7. 7.
    Duplicate Studies: Exclude studies that are duplicates or have been minimally modified for different publications, to ensure the uniqueness and quality of the analysis.
By setting these detailed inclusion and exclusion criteria, our literature review section can more accurately filter the highest quality literature, which will be most relevant to our research objectives. This helps enhance the depth and breadth of the study, while also ensuring the novelty and applicability of the research findings.
2.3 Step 3: Selection Process
In the process of selecting literature, we first comprehensively considered how the literature covers various factors affecting carbon credit prices, including market demand, policy changes, economic impacts, and technological advancements. Next, we evaluated various algorithms proposed in the literature, especially those applied to predicting carbon credit prices and optimizing carbon emission trading strategies, such as machine learning, artificial intelligence, and other statistical models. We also paid special attention to the performance evaluation, test results, and practical application cases of these algorithms. Furthermore, we explored research on how to predict corporate carbon emissions and its applications, especially those discussing the accuracy and feasibility of prediction models and their applications in corporate strategic planning, risk management, and compliance reporting. Finally, we focused on literature analyzing the impact of disclosing carbon credits on corporate financial performance, brand image, market competitiveness, and sustainable development strategies.
To ensure that the selected literature meets the standards of research quality, we developed the following Quality Assessment (QA) criteria:
  • QA1: Does the literature qualitatively and quantitatively explore the impact of non-disclosure of carbon emissions and carbon management information?
  • QA2: Does the literature conduct an in-depth analysis of factors influencing carbon credit prices, providing sufficient evidence and discussion?
  • QA3: Does the literature propose new methods or algorithms for predicting carbon credit prices and validate their effectiveness?
  • QA4: Does the literature propose new methods or algorithms for predicting corporate carbon emission and validate their effectiveness?
  • QA5: Does the literature clearly describe the research results and provide a thorough evaluation of models or algorithms?
In summary, we have collected a comprehensive set of research articles focusing on the different aspects of carbon credit systems, including factors influencing carbon prices, prediction algorithms for carbon credit prices and corporate emissions, and the impact of carbon emissions disclosure on organisations. This systematic approach ensures that our study incorporates high-quality literature, providing valuable insights into the dynamics of the carbon credit market and its intersection with financial technologies.
3 For FinTech: The Impact of Disclosing Carbon Emissions on organisations
As global climate change worsens, corporate carbon emission management and transparent information disclosure have become critical topics of concern. Carbon emissions are not only vital indicators of a company’s environmental footprint but also an essential measure of whether a company fulfills its sustainable development commitments. The disclosure of carbon emission information by organisations not only affects their market value but also has profound implications for the effectiveness of global carbon credit trading mechanisms. Given this context, this paper conducts an in-depth analysis of carbon emission disclosure practices in developed and developing countries, with a particular focus on the relationship between carbon emission disclosure and corporate market performance.
Studying carbon emission disclosure practices in these countries is of significant importance. First, differences in policies and market environments regarding carbon emission management and disclosure across countries can reveal the factors that drive or hinder organisations from disclosing carbon emission information. Second, an in-depth exploration of the impact of carbon emission disclosure can help understand how organisations can enhance market trust through transparent environmental responsibility practices and provide valuable insights for policymakers to further improve carbon credit trading mechanisms.
Recent studies in the United States suggest that firms with voluntary disclosure of carbon emissions experience higher valuations than their counterparts without voluntary disclosure. For example, a study conducted in the United States concluded that the market imposes stricter penalties on those who choose not to disclose relevant information, and that high carbon emissions have a negative impact on a company’s market value. [17].
Similar results have been replicated in the United Kingdom. A research in the United Kingdom, another developed country, shows that carbon emissions significantly negatively reduce corporate value, with markets typically reacting unfavorably to organisations with high carbon emissions. Even when organisations disclose more carbon emission information, the market response remains negative. This finding aligns with the results observed in the U.S. market [10]. Furthermore, the study have revealed a significant positive correlation between capital expenditure and carbon emission disclosure. Specifically, the more capital expenditure a company has, the more carbon emission information it discloses in its annual reports. This relationship is likely due to the fact that capital expenditure is typically associated with increased production activities, which in turn lead to higher carbon emissions [10].
Additionally, the study highlights that robust internal corporate governance, such as board size, board independence, and gender diversity enhances the level of carbon emission disclosure [10]. Other studies focusing on UK organisations have explored the relationship between carbon emission, disclosure, and financial performance. The findings indicate a negative correlation between carbon emission levels and corporate financial performance. Conversely, a significant positive correlation exists between carbon information disclosure and financial performance, indicating that organisations that voluntarily disclose carbon information experience significant improvements in financial performance [18, 19].
In the Australian market, disclosure policy plays a key role in influencing voluntary disclosure. The introduction of the National Greenhouse and Energy Reporting (NGER) Act in 2007 boosted voluntary disclosures of carbon emission, especially among larger, more visible firms. The research found that company size and carbon emission levels are key determinants of the extent of carbon emission disclosure. Larger organisations tend to disclose more carbon emission information, because they are more visible to the public and the regulators, thus having greater motivation to meet societal expectations through transparent disclosure practices. Furthermore, organisations with higher carbon emissions are more inclined to provide detailed disclosures, particularly in high-emission industries such as energy, transportation, materials, and utilities, where organisations face greater social and regulatory pressures, leading to more proactive and detailed disclosures [8].
The study also suggests that good corporate governance is crucial in promoting carbon emission disclosure. organisations with better governance structures (such as strong board independence and high governance scores) are more likely to disclose comprehensive carbon emission information, reflecting the positive role of good governance in enhancing corporate environmental transparency. Moreover, significant differences in disclosure practices exist across industries. organisations in high-emission industries, due to greater social and political pressures, generally disclose more detailed carbon emission information, while in low-emission industries, better disclosure practices are often observed in larger organisations or those with more robust governance structures [8].
Research in Japan further supports the aforementioned conclusions. Data analysis shows that reducing carbon emissions and disclosing carbon management activities have a positive impact on corporate market value. Specifically, the study reveals a significant negative correlation between corporate carbon emission levels and stock market value, while carbon management disclosure is positively correlated with stock market value. Additionally, the study finds that the positive correlation between carbon management disclosure and stock market value becomes more pronounced when carbon emission levels are higher. This suggests that reducing carbon emissions and actively disclosing related management activities are critical to a company’s sustainable development [20].
In contrast to developed countries, studies conducted in developing countries have different results. For instance, a study in Indonesia found no direct effect of carbon emissions disclosure on firm value. Investors in these markets tend to prioritize financial performance over environmental performance. However, the study also found that disclosing carbon emission information can improve corporate financial performance, as organisations that disclose such information can attract environmentally conscious consumers, thereby increasing revenue and profitability, which is consistent with observations in developed countries [11].
Further analysis suggests that effective government policy guidance is crucial for addressing climate change and enabling capital markets to hold organisations accountable for their carbon emissions. The implementation of mandatory disclosure policies, requiring organisations to publicly disclose their carbon emissions, is a key step in this process. Under such regulatory environments, the negative impact of carbon emissions on corporate market value becomes more significant, while voluntary disclosure of carbon emission information can partially mitigate this impact. It is through such mechanisms that the market can effectively penalize organisations with high carbon emissions and exert further pressure on those that choose not to disclose relevant information, which is why developed country governments generally adopt mandatory disclosure policies [11].
For example, a study in Malaysia found that corporate carbon emission levels and disclosure do not significantly impact market value, further indicating that in the absence of mandatory policies, organisations lack motivation to strengthen carbon management [21]. Similar findings were observed in South Korea, where the impact of carbon emissions and disclosure on corporate market value is limited, as the capital market in Korea places greater emphasis on corporate profitability rather than environmental impact. Therefore, effective government policy guidance is particularly important, requiring stricter regulations to direct market attention towards corporate carbon emission behaviors [22]. In contrast, a study in the UK demonstrated that the government’s mandatory disclosure requirements led to greater negative impacts on the market value of high-carbon-emission organisations, further confirming the critical role of policy in driving corporate carbon management and information disclosure [23].
In addition to exploring the impact of carbon emissions on corporate market value, our research also found that carbon emission information disclosure and actual emission levels significantly influence corporate loan interest rates. The study indicates that organisations that voluntarily disclose carbon emission information typically enjoy lower loan spreads, particularly in organisations with lower information transparency (such as those with incomplete financial information or insufficient credit ratings). For these organisations, voluntary disclosure of carbon emission information can significantly reduce loan spreads, thereby lowering borrowing costs. Furthermore, the study reveals a positive correlation between actual corporate carbon emission levels, particularly direct carbon emissions (i.e., Scope 1 emissions), and loan spreads. This means that higher carbon emission levels lead to higher loan spreads. Financial markets, especially banks, consider environmental performance, particularly carbon emission levels, when assessing corporate credit risk. High carbon emissions are viewed as potential environmental risks, leading to an increase in loan spreads [9].
This finding further underscores the importance of carbon emission information disclosure and carbon management. Transparent disclosure not only reduces negative impacts on the capital market but also brings tangible economic benefits to organisations by lowering loan costs.
Through a comprehensive analysis of carbon emission information disclosure research across multiple countries, we have drawn the following clear conclusions:
  1. 1.
    Impact of Carbon Emissions on Corporate Market Value: In studies conducted in developed countries such as the United States, the United Kingdom, Australia, and Japan, a significant negative correlation was found between corporate carbon emission levels and market value. In other words, organisations with high carbon emissions tend to experience a negative impact on their market value. However, through voluntary disclosure of carbon emission information, organisations can partially mitigate this negative impact. This suggests that transparent carbon emission disclosure plays an important buffering role in capital markets.
  2. 2.
    Impact of Carbon Emission Disclosure on Corporate Financial Performance: Research, particularly in the UK and Australia, shows a significant positive correlation between carbon emission disclosure and corporate financial performance. organisations that voluntarily disclose carbon information tend to experience improved financial performance, possibly due to market recognition and support for their environmental responsibility.
  3. 3.
    Impact of Carbon Emission Disclosure on Bank Loan Interest Rates: The study indicates that organisations that voluntarily disclose carbon emission information typically enjoy lower loan spreads, especially in organisations with lower information transparency. Conversely, higher carbon emission levels lead to increased loan spreads. This conclusion is applicable to both developed countries and some developing countries, indicating that financial institutions are increasingly considering corporate carbon emission behaviors when assessing risks.
  4. 4.
    The Key Role of Government Policy in Carbon Emission Disclosure: Studies in developing countries (such as Indonesia and Malaysia) have shown that due to the lack of mandatory disclosure policies, carbon emission information disclosure has little impact on corporate market value. This highlights the crucial role of government policy guidance. In contrast, in countries like the UK with mandatory disclosure policies, the market value of high-carbon-emission organisations is more negatively impacted, further proving the critical role of policy in driving corporate carbon management.
Fig. 1 and Fig. 2 summarize the factors influencing corporate carbon emission disclosure and the positive impacts of corporate carbon emission disclosure respectively.
Figure 1: Factors Influencing Carbon Emission Disclosure
Figure 2: Impact of carbon emissions disclosure
In conclusion, organisations that engage in transparent carbon emission information disclosure and proactive carbon management can not only maintain a competitive edge in the market, gaining trust from capital markets and financial institutions, but also enhance financial performance, comply with government regulations, and reduce environmental risks, thereby promoting sustainable development. Therefore, organisations should integrate carbon management into their core strategies, actively disclose emissions, and pursue emission reductions to effectively respond to market and policy challenges, creating greater value for both themselves and society.
4 For Engineering: Factors Influencing Carbon Credit Prices
Building on the previous conclusion, organisations that engage in transparent carbon emission disclosure and proactive carbon management not only maintain a competitive edge in the market and earn the trust of capital markets and financial institutions, but also improve financial performance, comply with government regulations, reduce environmental risks, and thereby promote sustainable development. Therefore, organisations should integrate carbon management into their core strategies, actively disclose emissions, and pursue emission reductions to effectively respond to market and policy challenges, thereby creating greater value for both themselves and society.
However, solely relying on proactive management and disclosure is far from sufficient. To further optimize carbon management strategies, organisations need to purchase carbon credits when prices are low to reduce operational costs and maximize the economic benefits of carbon management. Research has found that for high-emission industries, carbon management costs will constitute a significant portion of expenses. Moreover, if carbon management costs are too high, the company’s competitive position will be weakened [19]. Therefore, understanding and analyzing the various factors that influence carbon credit prices is crucial. Through qualitative analysis of these factors, organisations can preliminarily predict carbon price trends, allowing them to more accurately determine the optimal timing for purchasing credits. Additionally, these factors can serve as input variables for carbon price prediction algorithms, enhancing the accuracy and applicability of predictions.
Next, this study undertakes a systematic review of the primary factors influencing carbon credit prices, as identified in existing literature, with the aim of providing valuable insights to support corporate decision-making in the carbon market. Table 1 provides a summary of the key factors that influence carbon credit prices and shows the experimental methods used to derive these conclusions. Table 2 offers a detailed breakdown of these influence factors across different carbon market products. The subsequent discussion synthesizes these findings and explores their implications.
Table 1: Summary of factors influencing carbon credit prices
Factor Findings Methodology and Ref.
Coal Rising coal prices lead industries to seek alternative energy sources with fewer carbon emissions, like oil or renewable energy, reducing the demand for carbon credits and thus reducing carbon prices. Econometric Model [24]
Static and Dynamic Panel Models [25]
LASSO and SVAR Models [26]
Multiscale Analysis Model [27]
Vector Error Correction Model (VEC) and GARCH [28]
OLS, IV, VAR model [29]
Econometric Model [30]
SVAR model [31]
Rising coal prices are typically associated with increased demand for traditional energy sources, leading to higher carbon emissions and increased carbon prices. BP Neural Network Model [32]
SEM and BN [33]
Feasible Quasi Generalized Least Squares (FQGLS) [34]
Crude Oil Oil price increases are closely related to overall economic activity and have a strong positive impact on carbon prices. Econometric Model [24]
BP Neural Network Model [32]
SV-TVP-VAR Model [35]
SEM and BN [33]
MEEMD and Fast Fourier Transform [36]
TVP-VAR Model [37]
Feasible Quasi Generalized Least Squares (FQGLS) [34]
SVAR model [31]
Increases in crude oil prices raise production costs, leading organisations to reduce supply, which lowers carbon emissions and reduces the demand and price of carbon emission futures. Vector Error Correction (VEC) Model [38]
SV-TVP-VAR Model [35]
Macroeconomics Growth in macroeconomic activity typically drive increases in industrial production and raw material consumption, leading to higher demand for carbon emissions and pushing carbon prices up. Econometric Model [24]
Vector Error Correction (VEC) Model [38]
BP Neural Network Model [32]
MEMD, VAR, VEC Models [39]
EEMD (Ensemble Empirical Mode Decomposition) [40]
Vector Error Correction Model (VEC) and GARCH [28]
SVAR model [31]
Similar Products An increase in similar products’ prices often leads to a rise in carbon credit prices, indicating a substitution relationship between the two. Econometric Model [24]
SV-TVP-VAR Model [35]
ARMA-GARCH [41]
Stock Market Indices Stock market indices have a significant positive impact on the carbon market SEM and BN [33]
Static and Dynamic Panel Models [25]
LASSO and SVAR Models [26]
ARIMA-GARCH Model, Random Forest Algorithm [42]
Multiscale Analysis Model [27]
Euro Index The Euro Index had a negative impact on carbon prices before COVID-19, mainly due to uncertainties like Brexit MEEMD and Fast Fourier Transform [36]
Utility Index The utility index positively impacts carbon prices, reflecting increased power plant activity and associated carbon emissions. MEEMD and Fast Fourier Transform [36]
Interbank lending rates Interbank lending rates negatively correlate with carbon prices; rising rates weaken market liquidity and reduce demand for carbon allowances, leading to lower carbon prices. Bai-Perron Structural Break Test and Newey-West Regression Model [43]
Exchange rate Exchange rates negatively impact carbon prices, particularly in the short and medium term. DCC-GARCH, TVP-VAR and GA-BP [44]
MEMD, VAR, VEC Models [39]
COVID-19 Pandemic The COVID-19 pandemic significantly reduced carbon prices. Bai-Perron Structural Break Test and Newey-West Regression Model [43]
EEMD (Ensemble Empirical Mode Decomposition) [40]
Clean Energy Rising clean energy prices will drive up carbon prices, while the development of clean energy may sometimes suppress carbon prices. BP Neural Network Model [32]
SEM and BN [33]
MEEMD and Fast Fourier Transform [36]
Static and Dynamic Panel Models [25]
TVP-VAR Model [37]
OLS, IV, VAR model [29]
Feasible Quasi Generalized Least Squares (FQGLS) [34]
Econometric Model [30]
SVAR model [31]
Electricity price The impact of electricity prices positively influence the price of carbon credits. TVP- VAR Model [37]
Multiscale Analysis Model [27]
OLS, IV, VAR model [29]
Econometric Model [30]
End product prices The rise in prices of end products such as minerals, steel, and paper can drive up carbon credit prices, as the increase in end product prices leads to higher production, thereby increasing the demand for carbon emissions. OLS, IV, VAR model [29]
Air quality Environmental factors like air quality significantly impact carbon prices; higher air quality index (AQI) levels are associated with rising carbon prices. ARMA-GARCH [41]
VAR-VEC model, GARCH model [45]
Vector Error Correction Model (VEC) and GARCH [28]
Temperature Temperature has a significant impact on carbon allowance trading volume. Extreme weather conditions, such as very cold or hot temperatures, increase energy demand, thereby affecting the carbon allowance market and driving up carbon prices. Econometric Model [30]
Autoregressive Distributed Lag Model (ARDL) [46]
Structural Break Test, Autoregressive Distributed Lag Model (ARDL) [47]
Supply factor The supply of carbon credits also affects carbon prices; when the supply is large, carbon prices decrease. Econometric Model [30]
Degree of Carbon Regulation Short-term negative impact of carbon regulation policies on carbon prices is greater than long-term effects, but carbon prices become more sensitive to policy information as market mechanisms improve. SV-TVP-VAR Model [35]
Interest Rates Interest rates have a negative short-term impact on carbon prices SV-TVP-VAR Model [35]
Interest rates have a positive impact on carbon prices, especially in the medium to long term, as higher interest rates significantly reduce carbon prices by increasing production costs. DCC-GARCH, TVP-VAR and GA-BP [44]
Gold Prices Gold prices positively affect carbon prices, indicating a relationship between increased gold mining activities and higher carbon emissions. MEEMD and Fast Fourier Transform [36]
Population growth rate There is a positive correlation between population growth rate and carbon prices. A higher population growth rate will increase future carbon prices, as population growth leads to higher carbon emissions, necessitating higher carbon prices to address climate change. EZ Climate Model and Uncertainty Analysis [48]
Growth rate of per capita output The growth rate of per capita output has a greater impact on carbon prices. A higher per capita output growth rate will significantly increase future carbon prices. EZ Climate Model and Uncertainty Analysis [48]
A systematic review of existing literature reveals that carbon credit prices are influenced by a variety of factors, such as Coal and crude oil. These factors include not only fluctuations in traditional energy markets but also macroeconomic indicators, regulatory policies, market sentiment, and external shocks. The following provides a detailed summary of these factors and their mechanisms of impact.
Energy prices play a crucial role in determining carbon credit prices. Increases in coal and crude oil prices typically lead to higher carbon emissions, which in turn drive up carbon credit prices. However, as coal prices rise, industrial sectors may shift towards alternative energy sources with lower carbon emissions, such as natural gas or renewable energy, which could reduce the demand for carbon credits and subsequently lower prices. The impact of natural gas prices on carbon credits can be twofold. On one hand, higher natural gas prices may elevate carbon credit prices. On the other hand, the development of renewable energy may temper this upward trend [24, 32, 35, 33, 36, 25, 26, 37, 27]. The above conclusion has been confirmed in multiple studies. For example, [26] justifies the negative relationship between coal prices and carbon emissions using the LASSO and SVAR models.
Macroeconomic indicators and stock market performance also significantly influence carbon credit prices. Economic growth often acorganisations increased industrial production and energy consumption, thereby raising the demand for carbon emissions and driving up carbon credit prices. Moreover, stock market indices, such as the Euro Stoxx 50 and S&P 500, play a critical role in the carbon market, particularly during periods of economic prosperity. However, in certain markets, such as China, the relationship between stock market performance and carbon prices may be negative, indicating that market maturity and investor expectations influence price fluctuations [24, 38, 32, 33, 25, 26, 42, 39, 27, 40]. Numerous studies have also validated the relationship between economic indicators and carbon prices. For example, [38] justifies the positive relationship between macroeconomic development and carbon prices using the Vector Error Correction (VEC) Model.
Regulatory policies have a significant impact on the carbon market. In the short term, stringent carbon regulation policies may increase compliance costs for organisations, thereby depressing carbon prices. However, as market mechanisms improve, the sensitivity of carbon prices to policy information may increase, potentially leading to positive impacts in the long run. Additionally, the price linkage between international carbon markets is another key consideration, particularly the positive impact of European carbon markets on China’s carbon market, which has become more evident since 2020 [35, 25, 43, 37].
Financial factors such as exchange rates and interest rates also have complex effects on carbon credit prices. Exchange rate fluctuations generally have a negative impact on carbon prices in the short and medium term. For example, when the Renminbi depreciates, carbon prices may rise in the short term but decline in the long term. Rising interest rates typically suppress market liquidity, reducing demand for carbon credits and leading to lower prices. However, in the medium to long term, rising interest rates may increase production costs, reducing demand for carbon emissions and thereby pushing up carbon prices [43, 44, 39]. These conclusions have also been validated in several studies. For example, [43] justifies the negative relationship between interbank lending rates and carbon prices using the Bai-Perron Structural Break Test and Newey-West Regression Model.
Market sentiment and significant events, such as the COVID-19 pandemic, have profound effects on carbon credit price volatility. During the pandemic, global economic uncertainty and supply chain disruptions led to a sharp decline in carbon prices. However, the implementation of the European Union’s 750 billion euro green recovery plan helped stabilize the carbon market, leading to a rebound in carbon prices. Additionally, fluctuations in market sentiment may exhibit nonlinear relationships, adding complexity to price predictions. In such cases, nonlinear models like random forests have demonstrated superior accuracy in predicting price changes [36, 43, 44, 41, 24].
Air quality is a significant environmental factor that influences carbon prices. Higher Air Quality Index (AQI) levels are typically associated with rising carbon prices, as poorer air quality often indicates increased carbon emissions, thereby driving up carbon prices. Several studies have confirmed this relationship [41, 45, 28]. For instance, [41] validated this positive correlation using the ARMA-GARCH model.
Temperature has a notable impact on carbon allowance trading volumes. Extreme weather conditions, such as very cold or hot temperatures, increase energy demand, which in turn affects the carbon allowance market and drives up carbon prices. Multiple studies have validated this effect using various econometric models [30, 46, 47].
The supply of carbon credits also affects carbon prices. When the supply of carbon credits is large, prices tend to decrease due to market oversupply. [30] demonstrated the inverse relationship between supply levels and carbon prices using an econometric model.
Gold prices have a positive effect on carbon prices, as increased gold mining activities are typically associated with higher carbon emissions, which raises the demand for carbon credits. [36] used the MEEMD and Fast Fourier Transform models to demonstrate the positive relationship between gold prices and carbon prices.
There is a positive correlation between the population growth rate and carbon prices. A higher population growth rate is expected to increase future carbon prices, as population growth is often accompanied by higher carbon emissions. [48] used the EZ Climate Model and Uncertainty Analysis to demonstrate the impact of population growth on carbon prices.
The growth rate of per capita output significantly influences carbon prices. Higher per capita output growth rates will lead to substantial increases in future carbon prices, as this growth is typically accompanied by increased industrial activity and energy consumption. [48] validated the influence of per capita output growth on carbon prices using the EZ Climate Model and Uncertainty Analysis.
In addition to the factors mentioned above, other influences have also been shown to affect carbon prices. For example, [41] justifies the negative relationship between the prices of similar products using the ARMA-GARCH model. Electricity prices have also been confirmed by numerous studies to have a positive impact on carbon prices [37, 27, 29, 30]. For example, [30] used an Econometric Model to validate this effect.
In conclusion, carbon credit price fluctuations result from the interplay of various factors. Energy prices, macroeconomic activity, regulatory policies, financial market dynamics, market sentiment, and external events collectively shape the dynamics of the carbon credit market. Understanding the interactions among these factors is crucial for predicting carbon price trends and optimizing corporate carbon management strategies.
Table 2: Influence Factors on Different Carbon Market Products
Market Factor Reference
European Union Allowance (EUA) Black Energy (Coal, Oil) and Clean Energy (Natural Gas, Solar) BP Neural Network Model [32]
Gold Price, Euro Index and Utility Index MEMD and FFT [36]
Stock Market Indices (SP 500, CAC 40, DAX) and Carbon Emission Market (EUA, CER) SEM and BN [33]
Market Stability Reserve (MSR) N-A MEMD, VAR, VEC Models [39]
Energy Market and Stock Market LASSO and MSVAR Models [26]
COVID-19 Pandemic Bai-Perron Structural Break Test and Newey-West Regression Model [43]
China’s Carbon Market Degree of Carbon Regulation, Crude Oil Price, EUA Futures Price and Interest Rates SV-TVP-VAR Model [35]
Exchange Rate and Coal Price DCC-GARCH, TVP-VAR and GA-BP [44]
Industrial Production, Stock Market Index and Renewable Energy Generation Static and Dynamic Panel Models [25]
Industrial Structure, Policy Factors, Environmental Factors and External Market ARMA-GARCH [41]
Energy Prices, Economic Indicators, Macroeconomic Variables and Market Sentiment ARIMA(1,1,1)-GARCH(1,1) Model, Random Forest Algorithm [42]
Significant Short-term Events EEMD [40]
Table 2 summarizes the key factors influencing the European Union Allowance (EUA) market and China’s carbon market, along with the relevant research models and methodologies. This table clearly illustrates the differences in influencing factors across these markets.
In the EUA market, carbon prices are affected by energy prices, financial market indicators, policy factors, and external events. In terms of the energy market, price fluctuations of black energy (such as coal and crude oil) and clean energy (such as natural gas and solar) are major drivers of carbon prices. As traditional energy prices rise, organisations tend to shift toward low-carbon energy, thereby affecting the supply and demand of carbon credits [32]. Regarding financial market indicators, factors such as gold prices, the Euro Index, and the utility index also influence carbon prices [33, 36, 26]. These indicators reflect macroeconomic and financial market changes, and their connection to the energy market has been verified through models such as MEMD and FFT [36]. As for stock market and carbon market indices, stock market indices (such as SP 500, CAC 40, DAX) have a significant positive impact on the carbon market, as they reflect economic growth expectations, which in turn affect energy demand and carbon emissions [33, 26]. Additionally, the Market Stability Reserve (MSR) is a policy mechanism unique to the EUA market, designed to stabilize the market and prevent excessive carbon price fluctuations, with its impact analyzed through VAR and VEC models [39]. The COVID-19 pandemic, as an external event, caused short-term shocks to the carbon market, most notably the sharp decline in carbon prices in 2020, followed by a recovery driven by stimulus plans [43].
In contrast, China’s carbon market is influenced by more policy and structural factors. Carbon regulation intensity is a critical factor in China’s carbon market, as the implementation of carbon regulation policies has a direct short-term impact on carbon prices, particularly by increasing compliance costs for organisations and suppressing market demand [35]. Crude oil and coal prices also play a crucial role in China’s market, with coal price fluctuations being especially prominent, reflecting China’s reliance on coal in its energy structure [44]. In addition, exchange rates and interest rates significantly impact China’s carbon market, particularly in the short and medium term, as fluctuations in exchange rates and interest rates affect market liquidity and energy prices [44, 35]. Lastly, the growth of industrial production and renewable energy also has a significant influence on China’s carbon prices. These factors reflect the level of industrialization in China and its ongoing energy transition, with increases in industrial production and renewable energy generation being closely linked to carbon market prices [25, 40].
Table 1 and Table 2 together provide valuable reference tools for future researchers, helping them gain a comprehensive understanding of the key factors influencing carbon credit prices and the analytical methods used to examine them. Table 1 summarizes the critical factors affecting carbon credit prices and presents the experimental methods used to derive these conclusions, helping researchers identify broadly applicable variables and validation models for the carbon market. Table 2 further refines this by demonstrating how these influencing factors manifest across different carbon market products, enabling researchers to select appropriate analytical directions and models based on the characteristics of their target markets.
5 Algorithms for Predicting Carbon Credit Prices
After thoroughly reviewing the factors influencing carbon prices, it becomes clear that these factors are not only critical for understanding carbon market fluctuations but also serve as essential input variables for carbon price prediction algorithms. Effectively incorporating these factors into predictive models can significantly enhance the accuracy of carbon price forecasts, making the review of these algorithms even more meaningful and necessary. By accurately predicting carbon prices, organisations can strategically plan their carbon credit purchases when prices are lower, thereby better managing their costs. This not only helps organisations control overall expenditures and strengthen their competitiveness in the market but also encourages them to actively participate in carbon management and transparently disclose their carbon emissions. In doing so, organisations can gain greater trust and advantage in markets that increasingly prioritize ESG principles. The following sections will now review the latest and effective carbon price prediction algorithms.
Sun and Wang used a combined method of empirical mode decomposition [49] and least squares support vector machine to predict carbon prices. They proposed a hybrid model based on factor analysis, empirical mode decomposition, improved particle swarm optimization [50], and least squares support vector machine [51] to consider historical carbon prices and external factors influencing carbon prices. The study investigated three typical carbon markets in China, and the results showed that this hybrid model was more accurate in predicting carbon prices than other comparable models [52].
Nadirgil employed the Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) [53] technique to decompose the carbon price time series into multiple Intrinsic Mode Functions (IMFs), capturing various frequency characteristics within the data. Subsequently, Permutation Entropy (PE) was utilized to classify the IMFs, grouping IMFs with similar complexity into Cooperative Intrinsic Mode Functions (Co-IMFs) to reduce model computational complexity and the risk of overfitting. For high-frequency Co-IMFs, Variational Mode Decomposition (VMD) was further applied for secondary decomposition to mitigate the impact of high-frequency noise on model prediction. Ultimately, the study integrated multiple machine learning models, including Long Short-Term Memory (LSTM), Gated Recurrent Unit (GRU), Multilayer Perceptron (MLP), and Backpropagation Neural Network (BPNN), and optimized model parameters using Genetic Algorithm (GA) to construct 48 different hybrid prediction models. The effectiveness of the dual decomposition method (CEEMDAN-VMD-BPNN-GA) in addressing prediction complexity was ultimately demonstrated [54].
Hao and Tian proposed an innovative carbon price prediction framework. First, the framework applies ICEEMDAN [55] to decompose the carbon price time series, effectively removing noise while retaining key signal features. Then, feature selection is conducted using the Partial Autocorrelation Function (PACF) and Maximum Relevance Minimum Redundancy (mRMR) methods. Next, the parameters of the KELM model are optimized using the Multi-Objective Chaotic Sine-Cosine Algorithm (MOCSCA). Finally, a multi-step prediction strategy is designed using the recursive method, which not only accurately predicts short-term price fluctuations but also provides reliable support for long-term trend forecasting [56].
Li and Liu proposed a carbon price prediction model based on secondary decomposition and feature selection. First, the carbon price time series is subjected to multi-level decomposition using Discrete Wavelet Transform (DWT) [57] and ICEEMDAN, breaking it down into sub-sequences of different frequencies. Simultaneously, a three-stage feature selection method is employed to identify external influencing factors that are highly correlated with carbon prices. Subsequently, Support Vector Regression (SVR) [58] and Multilayer Perceptron (MLP) are used to predict the high-frequency and low-frequency sub-sequences, respectively, in order to capture short-term fluctuations and long-term trends. Finally, a weighted ensemble of the predictions from each sub-sequence is performed to obtain the overall carbon price prediction [59].
Wang et al. proposed a comprehensive carbon price prediction framework. Initially, the carbon price time series undergoes multi-level decomposition and denoising using Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) and Wavelet Transform (WT) to extract key features and reduce noise interference. Subsequently, Principal Component Analysis (PCA), Random Forest (RF), and Gradient Boosting Decision Tree (GBDT) are employed for feature selection, ensuring that the variables input into the model possess the strongest predictive capabilities. In the prediction phase, Bidirectional Gated Recurrent Unit (BIGRU) [60] is utilized for point prediction, leveraging bidirectional layers to capture the complex dynamic patterns within the time series. Following this, Gaussian Process Regression (GPR) [61] is applied for interval prediction, providing confidence intervals for the forecast results [62].
Liu and Shen proposed a carbon price prediction model based on Empirical Wavelet Transform (EWT) [63] and Gated Recurrent Unit (GRU), aiming to enhance prediction accuracy by integrating multiple techniques. First, the carbon price time series is decomposed using EWT, extracting components of different frequencies, including trend components, low-frequency components, and high-frequency components. Subsequently, Fuzzy C-Means (FCM) [64] clustering is employed to classify these components, further refining the signal characteristics. Then, the Partial Autocorrelation Function (PACF) is used to analyze the lag relationships of each component, determining the optimal lag order as input for the GRU model. Finally, the GRU model performs single-step predictions for each component, and the predicted results are aggregated to obtain the overall carbon price prediction [65].
Huang and He proposed a carbon price prediction model based on multi-scale joint prediction, integrating techniques such as Mean Optimization Empirical Mode Decomposition (MOEMD), Grey Relational Analysis (GRA), Factor Analysis (FA), Cooperative Kidney Algorithm (CKA) [66], and Extreme Learning Machine (ELM). First, MOEMD is used to decompose the carbon price data, extracting modal components at different frequencies. Then, GRA and FA are employed to select both structured and unstructured data highly correlated with carbon price fluctuations, including unstructured data like Baidu Search Index. Subsequently, CKA is utilized to optimize the parameters of the ELM model, enhancing prediction accuracy and stability. The experimental results demonstrate that the proposed MOEMD-CKA-ELM model significantly outperforms other comparative models in carbon price prediction [67].
Yang et al. proposed a carbon price prediction model based on a multi-factor combination, integrating Modified Ensemble Empirical Mode Decomposition (MEEMD), Long Short-Term Memory (LSTM) networks [68], and a rule-based machine reasoning system. First, MEEMD is employed to decompose the carbon price time series, extracting modal components of different frequencies, which provide clear signals for subsequent predictions [69]. Then, the LSTM model, with its strong capability for processing time series data, accurately predicts each modal component. To further enhance prediction accuracy, a rule-based machine reasoning system is introduced to automatically optimize the key parameters of the LSTM model, ensuring robust performance under various conditions. The experimental results demonstrate that the proposed MEEMD-LSTM model outperforms traditional models in carbon price prediction [70].
Lu et al. proposed a carbon price and trading volume prediction framework based on multi-model comparison, incorporating the Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) technique to evaluate the performance of six different machine learning models. The machine learning models used include XGBoost [71], Random Forest (RF) [72], Grey Wolf Optimizer-based Nonlinear Extension of the Arps Decline Model (GWO-KNEA) [73], Particle Swarm Optimization-based Support Vector Machine (PSO-SVM) [74], Simulated Annealing and Fruit Fly Optimization Algorithm-based Support Vector Machine (SA-FFOA-SVM), and Radial Basis Function Neural Network (RBFNN) [75]. Through experiments conducted on data from eight carbon trading markets in China, the authors compared the prediction accuracy and stability of each model. The study found that the CEEMDAN-GWO-KNEA and CEEMDAN-RBFNN models performed exceptionally well across multiple datasets [76].
Tian et al. proposed a novel hybrid model based on feature selection and multi-objective optimization algorithms for carbon price prediction. The study used Weighted Regularized Extreme Learning Machine (WRELM) as the artificial intelligence algorithm to improve the accuracy and stability of carbon price prediction [77]. The model overcame the limitations of statistical models in handling carbon price prediction and optimized the model using the Multi-Objective Grasshopper Optimization Algorithm (MOGOA), achieving excellent prediction results [78]. This research provides important references for the rational and stable development of the carbon market [79].
Yang et al. proposed a carbon price prediction model based on Modified Ensemble Empirical Mode Decomposition (MEEMD), Long Short-Term Memory (LSTM) networks, and an Improved Whale Optimization Algorithm (IWOA) [80]. First, MEEMD is used to decompose the carbon price time series into several Intrinsic Mode Functions (IMFs) and a residual component, reducing noise and mode mixing issues [81]. Subsequently, the LSTM model is employed to predict the future values of each IMF and the residual component, capturing long-term dependencies in the time series [68]. To further enhance prediction performance, IWOA is used to optimize the initial weights and biases of the LSTM model [82].
Zhao et al. proposed a carbon price prediction model that combines the Hodrick-Prescott (HP) filter, an improved Grey Model (Variable Grey Model, VGM), and Extreme Learning Machine (ELM) [83]. First, the HP filter is used to decompose the carbon price data and its influencing factors into long-term trends and short-term fluctuations, which are modeled separately. The VGM model is then employed to predict the long-term trends, while the ELM handles the short-term fluctuations. By combining the predictions from these two components, the model effectively enhances the accuracy of carbon price prediction [84].
Wang et al. proposed a carbon price interval prediction model based on Lower Upper Bound Estimation (LUBE), Asymmetric Multi-Objective Evolutionary Algorithm (MOEA), and Long Short-Term Memory (LSTM) networks [68]. First, the carbon price time series is decomposed using Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) to reduce noise and extract key features. Next, Symbol Transfer Entropy (STE) is employed for causal inference and feature selection to ensure the effectiveness of the input variables [85]. Then, based on the prediction points generated by LSTM, the LUBE model uses MOEA to optimize parameters and generate asymmetric prediction intervals [86].
Deng et al. proposed a carbon price trend prediction model that integrates several advanced techniques. The model employs Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) to decompose carbon price time series into Intrinsic Mode Functions (IMFs) and residuals, allowing the extraction of significant features. Following this, a Two-Stage Feature Selection (TFS) method is used to identify the most relevant features from these IMFs, which are then fed into a Light Gradient Boosting Machine (LightGBM) model [87]. The hyperparameters of the LightGBM model are optimized using Bayesian Optimization Algorithm (BOA) to enhance its predictive performance [88]. Finally, SHapley Additive exPlanations (SHAP) is applied to interpret the model’s predictions [89], providing insights into the contributions of various features to the predicted carbon price trends [90].
Shi et al. proposed a deep learning framework combining Convolutional Neural Networks (CNN) and Long Short-Term Memory (LSTM) networks to predict carbon prices with high accuracy and robustness. The model first uses CNN layers to extract spatial features from the carbon price time series data, which are then passed to LSTM layers to capture temporal dependencies [91]. This hybrid CNN-LSTM model is designed to address the limitations of traditional models in capturing complex nonlinear patterns in carbon markets [92]. The study demonstrates that the CNN-LSTM model significantly outperforms single CNN or LSTM models in both prediction accuracy and robustness, particularly in handling anomalies and outliers in the data [93].
Zhang et al. proposed a hybrid model for multi-step carbon price prediction, integrating Extremely Randomized Trees (ET), Multivariate Variational Mode Decomposition (MVMD), and Long Short-Term Memory (LSTM) networks. The model employs ET to select the most influential external factors affecting carbon price fluctuations, simplifying the model’s complexity and enhancing prediction accuracy. [94] MVMD is then used to decompose the carbon price and selected external variables into multiple modes, allowing for the extraction of stable frequency components that are crucial for accurate predictions. Each mode is subsequently modeled using LSTM to capture long-term dependencies in the time series, with the final prediction obtained by aggregating the individual predictions [95].
Zhang et al. proposed an enhanced decomposition-ensemble model for both point and probabilistic carbon price forecasting. The model integrates Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN), Fuzzy Entropy (FE) [96], Variational Mode Decomposition (VMD) [97], Long Short-Term Memory (LSTM) networks, Gated Recurrent Unit (GRU) networks, Cuckoo Search Algorithm (CS) [98], and Gaussian Process Regression (GPR). Initially, CEEMDAN is employed to decompose the carbon price series into Intrinsic Mode Functions (IMFs) and residuals, followed by FE to quantify the complexity of these IMFs. High-complexity IMFs undergo further decomposition using VMD. LSTM and GRU models then predict each decomposed component, with CS optimizing their weights. Finally, GPR is used to generate prediction intervals, offering probabilistic insights [99].
Hong et al. proposed a hybrid carbon price forecasting framework that integrates a Deep Augmented Frequency Enhanced Decomposed Transformer (DA-FEDformer) with a Multimodel Optimization Piecewise Error Correction System [100]. The DA-FEDformer model enhances frequency domain information extraction using a multi-layer perceptron integrated into the encoder-decoder layers, combined with an Improved Kernel Mean Square Error (IKMSE) loss function and the Lion Optimizer. This model generates initial carbon price predictions, which are then refined by the error correction system [55]. The system employs Improved Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (ICEEMDAN) to decompose the prediction error sequence and applies five different models to correct these errors [101].
Table 3: Summary of Price Prediction Methods and Techniques
Ref. Price prediction method Optimization method Data decomposition technique Multi-factor
[52] LS-SVM Improved Particle Swarm Optimization (IPSO) EMD Y
[54] LSTM, GRU, MLP, BPNN Genetic Algorithm (GA) CEEMDAN+Permutation Entropy (PE)+Variational Mode Decomposition (VMD) N
[56] KELM Multi-Objective Chaotic Sine Cosine Algorithm (MOCSCA) ICEEMDAN Y
[59] SVR (high-frequency subsequences) + MLP (low-frequency subsequences) - ICEEMDAN+Discrete Wavelet Transform (DWT) Y
[62] BIGRU point prediction + GPR interval prediction - CEEMDAN+Wavelet Transform (WT) Y
[65] GRU - Empirical Wavelet Transform (EWT) N
[67] Extreme Learning Machine, ELM Cooperative Kidney Algorithm (CKA) Mean value Optimization Empirical Mode Decomposition, MOEMD Y
[70] LSTM - MEEMD Y
[76] XGBoost, RF, KNEA, SVM, RBFNN GWO, PSO, SA, FFOA CEEMDAN N
[79] Weighted Regularized Extreme Learning Machine, WRELM Multi-Objective Grasshopper Optimization Algorithm, MOGOA ICEEMDAN Y
[82] LSTM IWOA MEEMD N
[84] VGM (long-term trend), ELM (short-term trend) PSO Hodrick-Prescott, HP filter Y
[86] LSTM, Lower Upper Bound Estimation (LUBE) Asymmetric Multi-Objective Evolutionary Algorithm, MOEA CEEMDAN Y
[90] Light Gradient Boosting Machine (LightGBM) Bayesian Optimization Algorithm, BOA CEEMDAN, Two-Stage Feature Selection, TFS Y
[93] LSTM - Convolutional Neural Network (CNN) N
[95] LSTM - Multivariate Variational Mode Decomposition, MVMD Y
[99] LSTM, GRU Cuckoo Search Algorithm, CS CEEMDAN, Fuzzy Entropy, FE, Variational Mode Decomposition (VMD) Y
[101] Deep Augmented Frequency Enhanced Decomposed Transformer (DA-FEDformer), Multimodel Optimization Piecewise Error Correction System - ICEEMDAN N
Summarizing the research progress in carbon credit price prediction over the past few years, we can observe a gradual evolution and enhancement of the techniques employed. Early studies primarily relied on single models, such as the Hybrid Neural Fuzzy Controller (PATSOS), Adaptive Neuro-Fuzzy Inference System (ANFIS), and Artificial Neural Networks (ANN). These models played a crucial role in the initial stages of carbon price prediction [102]. However, as data complexity increased, researchers began introducing data decomposition techniques, such as Empirical Mode Decomposition (EMD) and its improved versions. These techniques allow time series data to be decomposed into more regular components, making it easier for machine learning algorithms to capture underlying patterns in the data. Additionally, optimization algorithms have been widely applied to enhance the overall performance of the models.
As research has advanced, the focus has gradually shifted away from relying solely on carbon prices as the single input variable, with other factors influencing carbon prices also being incorporated into the models. For instance, energy prices, macroeconomic indicators, and environmental factors have been included, and these external variables are effectively selected using feature selection methods to ensure the quality and relevance of the model inputs.
To further improve prediction accuracy, recent research has concentrated on more refined data decomposition and optimized multi-model combinations. On the one hand, after initially decomposing the data, researchers have employed more precise secondary decomposition techniques to extract potential patterns, enabling the models to learn and predict more accurately. On the other hand, the combination of different prediction models and optimization algorithms has been extensively explored to identify the most suitable hybrid models for different carbon markets. Moreover, some studies have decomposed raw data into components with different frequencies and applied different prediction algorithms to each frequency component, ultimately integrating the predictions of each sub-model through weighted ensemble methods to enhance overall prediction performance.
In addition to point predictions, an increasing number of studies have also introduced interval prediction techniques to provide a prediction range for carbon prices. This approach not only improves the reliability of predictions but also better addresses the uncertainties present in the market, offering more comprehensive predictive insights.
6 Prediction of Enterprise Carbon Emissions
Corporate carbon emission prediction algorithms hold significant importance in the current global context of combating climate change. Beyond assisting organisations in better managing their carbon emissions, these predictive methods are also crucial in meeting the needs of investors and governments. Financial institutions are increasingly focused on whether organisations adhere to ESG principles and prefer to invest in those that can create long-term social value. Accurate carbon emission predictions help enhance corporate transparency, reduce information asymmetry, and strengthen market trust. Moreover, as governments impose stricter requirements on corporate carbon emission disclosures, precise carbon emission predictions can assist organisations in better complying with regulations, supporting government efforts to achieve emission reduction targets, and effectively evaluating whether organisations are truthfully disclosing and adhering to carbon emission rules. The following section will review various carbon emission prediction methods.
6.1 Empirical model:
In the context of predicting corporate-level carbon emissions, the study by Goldhammer et al. presents an innovative approach that estimates corporate carbon footprints (CCFs) using externally available data. They employ a regression analysis model that primarily utilizes five key variables as input data: firm size (measured by turnover or number of employees), vertical integration, capital intensity, production centrality, and the carbon intensity of the national energy mix. The model was trained on actual CCF data from 93 European organisations across the chemical, construction engineering, and industrial machinery sectors. The study demonstrates that this method can accurately estimate corporate carbon emissions without relying on internal data, making it particularly suitable for scenarios requiring rapid, external assessments [103].
Shakdwipee and Lee explores the application of two primary technical methods. The first is based on the Economic Input-Output Life Cycle Assessment (EIO-LCA) model, which utilizes macroeconomic data, such as industry GDP and carbon emission intensity, to estimate corporate carbon emissions. The second method is the estimation model developed by MSCI ESG Research, which combines historical company disclosures with industry-specific intensity to predict carbon emissions for years without disclosed data. The research used the 2013 carbon emission data from 277 organisations in the MSCI ACWI IMI index as training data and evaluated the models’ prediction accuracy by comparing them with the organisations’ actual disclosed data. The results indicate that the MSCI ESG Research model outperforms the EIO-LCA model in estimation accuracy, particularly in handling company-specific carbon emission intensities [104].
The study by Griffin et al. provides an effective technical approach for estimating corporate greenhouse gas (GHG) emissions using external data. This research employs the Ohlson valuation model [105] and combines it with reported GHG emission data to develop a GHG emission estimation model through regression analysis. This model utilizes a range of company-specific variables as input data, including revenue, capital expenditures, and asset composition, to predict the emissions of organisations that have not disclosed their GHG data [106].
6.2 Machine learning method:
Zhou et al. developed a carbon risk prediction model based on Support Vector Machines (SVM) to accurately predict carbon emissions for heavily polluting industrial enterprises in China. The study first established a carbon risk assessment system comprising 45 initial indicators. After performing principal component analysis and independent analysis, 18 core indicators were selected. The SVM algorithm was used for model training, and the penalty parameter C and Gaussian kernel parameter gamma were optimized through grid search and cross-validation to handle the nonlinear and high-dimensional characteristics of the data [107].
Quyen et al. developed a novel approach for predicting corporate-level carbon emissions, specifically targeting organisations that do not publicly disclose their emissions data. The framework proposed a Meta-Elastic Net learning model to aggregate predictions from multiple base models, including Ordinary Least Squares (OLS), Elastic Net, Neural Networks, K-Nearest Neighbors (KNN), Random Forest (RF), and Extreme Gradient Boosting (XGBoost). The input training data comprised global carbon emissions data from 2,289 organisations between 2005 and 2017, totaling 13,435 annual observations. The data included Scope 1, Scope 2, and Scope 3 emissions, as well as various corporate factors such as operational scale, business models, technological advancements, and energy consumption. The model’s performance was validated through extensive cross-validation and robustness tests, demonstrating significant improvements in prediction accuracy compared to traditional regression models [108].
Li et al. proposed an advanced carbon emission prediction model tailored for enterprise-level applications, integrating multi-dimensional feature extraction with an optimized Radial Basis Function (RBF) neural network. The model utilizes Principal Component Analysis (PCA) [109] combined with kernel functions to extract essential features from a diverse set of input data, which includes production metrics, fuel consumption, electricity and heat energy purchases, and other relevant economic indicators. This approach effectively reduces data complexity while preserving key predictive information. The extracted features are then input into an RBF neural network, whose parameters are finely tuned using a hybrid optimization method that combines the Drosophila Optimization Algorithm (DOA) and Differential Evolution (DE). This dual optimization strategy significantly enhances the accuracy and robustness of the emission predictions [110].
Chen et al. proposed an efficient Fast Transfer-Based Adaptive Recurrent Neural Network (SEER) model aimed at addressing the high computational costs associated with real-time predictions in existing deep learning models. The SEER model enhances prediction accuracy and training efficiency by integrating data preprocessing, feature extraction, attention mechanisms, transfer learning, and adaptive learning rate mechanisms. Using carbon emission and air quality data from Shanghai and Beijing as inputs, the model effectively identifies key data points and optimizes generalization capabilities, demonstrating its accuracy and applicability in corporate carbon emission predictions [111].
Tang et al. proposed CarbonNet, a novel approach for enterprise-level carbon emission prediction, aimed at overcoming the limitations of traditional macro-level forecasting models. The paper first constructs a large-scale dataset encompassing 3346 organisations over 31 years, integrating carbon emission data with corporate financial and operational metrics. To enhance prediction accuracy, the authors employed Factor Analysis for feature extraction, reducing the impact of irrelevant variables, thereby building a more generalizable carbon emission prediction model. Subsequently, the paper utilized the XGBoost algorithm for large-scale data mining, iteratively optimizing prediction errors, resulting in a median absolute relative error (MARE) of 0.25, which represents a 22% improvement in prediction accuracy compared to existing methods [112].
Shou developed a corporate carbon performance evaluation framework using Support Vector Machine (SVM) to predict carbon emissions at the enterprise level. The framework, named Enterprise Carbon Performance Index System (ECPIS), integrates SVM with real-time data collection and monitoring to assess and forecast carbon emissions, enabling organisations to make data-driven decisions to reduce their carbon footprint. The input training data for the SVM model includes historical carbon emissions data, energy consumption metrics, operational data such as production rates and operating costs, as well as environmental data from air quality sensors and greenhouse gas analyzers. The system demonstrates a 97.5% accuracy in evaluating corporate carbon performance, outperforming other models like K-Nearest Neighbors (KNN), Random Forest (RF), Decision Trees (DT), Naive Bayes (NB), and Logistic Regression (LR) [113].
Zhao et al. proposed an innovative approach for predicting carbon emissions at the enterprise level, specifically targeting coal combustion processes in power generation organisations. This model utilizes Rough Set Theory to preprocess and extract key features from incomplete and imprecise carbon emission data, thereby enhancing the reliability of the data for subsequent modeling. The model then integrates a Grey System GM(1,1) model with a Support Vector Machine (SVM) to predict carbon emissions. The Grey model generates initial predictions based on historical data, which are further refined by the SVM to account for nonlinear relationships and improve accuracy. The input data for the model includes coal consumption rates, corresponding carbon emissions, and other related variables from a large coal-fired power company, covering multiple plants. The combination of the Grey model and SVM results in a hybrid prediction framework that significantly reduces prediction errors compared to traditional methods [114].
Shi et al. developed a carbon emission prediction model tailored for thermal power enterprises, utilizing data from 17 power plants in Gansu Province, China. The study introduced a hybrid approach that combines Factor Analysis (FA) with multiple machine learning models, including Support Vector Regression (SVR) and a Genetic Algorithm-optimized Backpropagation Neural Network (GA-BP). The model inputs include critical variables such as energy consumption, electricity output, fuel type, and carbon content per unit of heat, derived from the 2021 greenhouse gas emissions reports and carbon verification documents of the power plants. Factor Analysis was employed to reduce data dimensionality by extracting key factors—such as energy consumption and output, energy quality, and energy efficiency—which were then used as inputs for the prediction models. Among the models evaluated, the GA-BP neural network demonstrated superior performance in predicting CO2 emissions, showcasing lower error rates and higher accuracy [115].
Pan et al. introduced a real-time estimation method for the cement industry, utilizing a CNN-BLSTM (Convolutional Neural Network-Bidirectional Long Short-Term Memory) model. This method accurately estimates both direct and indirect carbon emissions by integrating CNN for feature extraction from material and energy consumption data, and BLSTM [116] for capturing temporal dependencies. Additionally, the study employs a Marginal Carbon Emission Factor (MCEF) to assess indirect emissions. The model was trained using 30 days of operational data from a cement corporation, demonstrating superior accuracy compared to traditional methods [117].
Figure 3: Evolutionary Trends in Corporate Carbon Emission Prediction Algorithms
After a thorough review of corporate carbon emission prediction algorithms, several conclusions can be drawn. The research in this field is increasingly moving towards more diverse data inputs, refined model processing, and industry-specific customization. By incorporating complex feature extraction techniques and multidimensional data inputs, researchers have enabled prediction models to more accurately capture the various factors influencing corporate carbon emissions. Moreover, the widespread application of machine learning and deep learning algorithms, along with the exploration of hybrid models and optimization algorithms, has significantly enhanced the accuracy and robustness of these prediction models. As the demand for real-time predictions grows, especially in high-emission industries, researchers have developed customized prediction models tailored to specific sectors, improving the effectiveness of these models, and Fig. 3 reflects the trends in the evolution of these technologies. These advancements not only assist organisations in better managing their carbon emissions but also provide investors and governments with more transparent and reliable carbon emission information, thereby supporting global efforts to achieve climate goals.
7 Prospects of Carbon Credits in the Financial and FinTech Fields
The existing literature has extensively studied the impact of carbon emissions and carbon information disclosure on corporate market value, financial performance, and loan interest rates. However, the role of carbon credits in corporate valuation, particularly in the financial and FinTech fields, remains an area with significant potential for further exploration.
Firstly, the application of carbon credits holds substantial importance for quantifying corporate market value. By predicting carbon emissions and carbon credit prices, organisations can estimate future costs associated with offsetting emissions. This forecasting not only helps organisations make more informed decisions regarding the timing and volume of carbon credit transactions but also allows for the development of internally trained algorithmic models that assist in decision-making and further manage operational costs. In this process, organisations can effectively reduce carbon emission costs and voluntarily disclose carbon information, thereby enhancing market trust, lowering financing costs, and ultimately increasing market value. This research direction lays a strong foundation for future studies to quantify the long-term impact of carbon management on corporate market value.
Secondly, the impact of carbon credits on corporate valuation warrants further investigation. By understanding the future costs incurred from purchasing carbon credits or the revenues generated from selling them, financial analysts can use the Discounted Cash Flow (DCF) method to discount these future costs/revenues to their present value. This present value directly affects a company’s market value, helping managers gain a clearer understanding of how carbon emissions influence financial performance and market valuation. This direction offers promising applications not only in traditional finance but also in the FinTech sector.
Finally, carbon risk ratings are also an important area of research. With the advancement of carbon emission prediction algorithms and price forecasting models, corporate carbon risks can be assessed with greater accuracy. These risk ratings can provide a basis for corporate valuation and help investors better understand a company’s environmental risks and future financial performance. By incorporating carbon emissions and carbon credit costs into multifactor models, predictions of corporate valuation will become more comprehensive and precise.
In conclusion, the prospects for carbon credits in the financial and FinTech fields are vast. Future research can further explore how carbon emission predictions, price forecasts, and risk ratings can provide more accurate quantitative tools for corporate valuation and market performance. This will not only help organisations optimize their carbon management strategies and enhance market competitiveness but also promote the development of the global carbon credit market, contributing to the achievement of sustainable development goals.
8 Literature Review Comparison
Currently, there is a very limited number of literature reviews on the topic of carbon credits. Below is a list of the most recent reviews concerning carbon credits and carbon management.
Trouwloon reviewed the use of carbon credits in corporate climate statements, focusing on the classification of corporate climate statements, quality issues related to carbon credits, and the aspects of risk and governance within these statements. By examining the role of carbon credits in corporate climate statements, this study provides a deeper understanding of corporate climate declarations, emphasizing the responsibilities and challenges organisations face in climate action, as well as how transparency and sustainability can be improved [13].
Lokuge and Anders conducted a comprehensive literature review on the carbon credit system in Canada’s agricultural sector, with a particular focus on Alberta. They discussed the impact of carbon tax policies on various industries, the research on global carbon credit markets, and the carbon credit systems in the agricultural sectors of developed countries, with special emphasis on Alberta and the western regions of Canada [14].
Alshatri and Hussain conducted a systematic literature review identifying the external factors influencing carbon credit prices and categorized them into energy market factors, environmental variables, macroeconomic indicators, social factors, and online search indices. They emphasized the complexity and multidimensionality of these factors [2].
Jin et al. systematically reviewed 147 carbon emission prediction models (CEPM) conducted between 2011 and August 2022, with a focus on model classification, optimization techniques, and influencing factors. However, this review primarily focuses on the prediction of carbon emissions at the national, industrial, or city level, rather than at the corporate level, thus providing limited insight for investors and governments to understand the specifics of individual organisations [118].
Wang et al. reviewed the impact of carbon risk on corporate capital costs and market value, highlighting the negative effects of high carbon emissions on corporate financial performance. However, Wang et al.’s review mainly concentrated on the direct impact of carbon emissions on organisations and did not fully explore the role of carbon emission disclosure in corporate management and market response [119].
In contrast, this study, through a systematic literature review (SLR), not only analyzes the impact of carbon emissions on organisations but also delves into the role of carbon emission disclosure and the latest technologies in carbon credit pricing and corporate carbon emission prediction. Thus, it provides more comprehensive guidance for organisations, financial investors, and governments.
Specifically, this study surpasses existing literature in the following aspects:
  1. 1.
    A more comprehensive analytical perspective: This study not only investigates the direct impact of carbon emissions and carbon credits on corporate market value but also focuses on the transparency issues in corporate carbon emission management and disclosure. This transparency not only affects corporate market performance but also has significant implications for corporate financing costs and investor trust.
  2. 2.
    Filling the research gap in carbon credit pricing and corporate carbon emission prediction technologies: Although existing literature has addressed the application of carbon credits in specific fields (such as agriculture or construction), this study systematically reviews these technologies in the financial and fintech sectors, revealing their importance to overall corporate management and market competitiveness.
  3. 3.
    Stronger practical guidance: This study not only analyzes the influencing factors of carbon credit prices but also evaluates the applicability and limitations of current carbon credit price prediction algorithms, helping organisations make more informed decisions in the carbon market and reduce management costs. Additionally, the study explores the latest developments in corporate carbon emission prediction technologies, assisting investors and regulators in more accurately assessing corporate environmental performance and reducing information asymmetry.
  4. 4.
    Implications for future research: By revealing the long-term impact of carbon credits and carbon emission management on corporate market value, this study provides a solid foundation for future research, particularly in quantifying the long-term effects of carbon management on corporate financial and market performance.
In conclusion, this study, through a broader and more in-depth literature review, addresses the shortcomings of existing literature on carbon credits and carbon emission management, offering more comprehensive theoretical and practical guidance for organisations, investors, and governments in addressing the challenges posed by climate change.
9 Conclusion
The study on carbon disclosure suggests that organisations should actively manage and accurately disclose their carbon emissions to mitigate the negative impact on market value and financial performance. By maintaining transparency, organisations can enhance market trust, reduce financing costs, and ultimately increase their market value.
Furthermore, research indicates that understanding the factors influencing carbon prices and accurately predicting carbon prices are crucial for corporate cost management. Accurate carbon price forecasts enable organisations to better plan the timing and volume of carbon credit transactions, thereby optimizing operational cost management.
Carbon emission predictions play a significant role in increasing corporate transparency. They not only help governments and financial investors gain a clearer understanding of actual corporate emissions but also provide a reliable basis for assessing environmental risks and future financial performance.
Finally, integrating carbon price predictions with carbon emission forecasts allows organisations to manage their carbon emission costs more comprehensively. The combination of these two techniques lays the foundation for future quantitative research, providing more precise tools for corporate valuation and market performance, helping organisations achieve sustainable development goals while enhancing market competitiveness.
In this way, these advancements not only assist organisations in optimizing their carbon management strategies but also promote the development of the global carbon credit market, contributing to the achievement of sustainable development goals.